OpenStudy (jacky90):

Two cyclists leave towns 180mi apart at the same time and travel toward each other. One cyclist travels 8 mi/h faster than the other. If they meet in 5 hours, what is the rate of each cyclist?

OpenStudy (anonymous):

Refer to the attachment.

OpenStudy (zpupster):

distance = 180mi rate of slower cyclist = x rate of faster cyclist 8+x distance = speed x travel time so, 5x + 5(x+8) = 180 5x + 5x + 40=180 10x+40 -(40)=180 -(40) 10x=140 your x is the rate of slower cyclist remember to do a check

OpenStudy (jacky90):

Thank you so much fellas :D